Find the areas of a parallelogram with sides of 12 ft and 8 ft and one angle of 45 degrees.
Answer by Theo(2973)

the area of a parallelogram is equal to the base times the height.

looks like you just to find the height.

since one angle of the parallelogram is equal to 45, then we'll use that angle.

if your parallelogram is abcd with a on the top left and b on the top right and c on the bottom right and d on the bottom left, then we let our angle = abc which means the bottom right corner of the parallelogram will form the 45 degree angle.

we let the left and right side of the parallelogram equal 8 feet each.

we let the bottom and top side of the parallelogram equal 12 feet each.

12 feet is the length of the base of the parallelogram.

the height will equal sin(45) * 8 = 5.65685425

area = 5.65685425 * 12 = 67.88225099 square inches.